OpenStudy (anonymous):

Verify the following identity: Tan2 x/ 1+ tan2 x = sin2 x

OpenStudy (anonymous):

Someone answer me please?!?:)

hero (hero):

\[\frac{\tan^2x}{1 + \tan^2x} = \sin^2x\]

OpenStudy (anonymous):

Yes that is the question but I need to find the identity and how I did it.. Make one side equal the other .. Idk how to do that with this one problem

hero (hero):

Well, basically since the right side is just a single term, we'll need to manipulate the left side so that it is also one term. The easiest way to do that is to turn \(1 + \tan^2x\) into an expression that is just one term. We can do that by expressing it like so:

hero (hero):

As you know: \(\tan^2x = \frac{\sin^2x}{\cos^2x}\) So we can re-write the left hand side as: \[\dfrac{\tan^2x}{1 + \dfrac{\sin^2x}{\cos^2x}}\]

hero (hero):

Now we can convert 1 to \(\dfrac{\cos^2x}{\cos^2x}\): \[\dfrac{\tan^2x}{\dfrac{\cos^2x}{\cos^2x} + \dfrac{\sin^2x}{\cos^2x}}\]

hero (hero):

Then combine fractions in the denominator to get: \[\dfrac{\tan^2x}{\dfrac{\cos^2x +\sin^2x}{\cos^2x}}\] Notice that \(\cos^2x + \sin^2x = 1\) so \[\dfrac{\tan^2x}{\dfrac{1}{\cos^2x}}\]

hero (hero):

Next express the fraction in the following form: \(\tan^2x \div \dfrac{1}{\cos^2x}\) And use the reciprocal rule to express it as: \(\tan^2x \times\cos^2x\) Express \(\tan^2x\) in explicit form: \(\dfrac{\sin^2x}{\cos^2x} \times \cos^2x\) And I'll leave the rest to you

OpenStudy (anonymous):

Ok thanks a lot!!!

hero (hero):

You're welcome.